OpenStudy (anonymous):

Evaluate, if possible: 4√2401

jimthompson5910 (jim_thompson5910):

\[\large \sqrt[4]{2401}\] \[\large \sqrt[4]{7^4}\] \[\large 7\] So \[\large \sqrt[4]{2401}=7\]

OpenStudy (anonymous):

\[\sqrt{2401}\] = 49 49 times 4 =196